OpenStudy (anonymous):

Megan and Sam together take 12 minutes to wash a car. Washing the car by herself Megan took 21 minutes. How long would it take Sam to wash the car?

OpenStudy (anonymous):

\[\frac1x + \frac{1}{21} = \frac1{12}\]Solve for x.

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

I'm guessing no??

OpenStudy (anonymous):

Woah sorry umm k gimme a sec

OpenStudy (anonymous):

x = -9?

OpenStudy (anonymous):

First step: Find a common denominator and set up a proportion.\[\frac{x + 21}{21x} = \frac{1}{12}\]Step 2: Cross Multiply 12(x + 21) = 21x Step 3: Solve for x 12x + 252 = 21x 9x = 252 x = ?